I can not get this one thanks for the help
The sum of a number and –11 is 4. Find the number

Answer by rapaljer(4671)

Let x = the number and write an equation:
x+(-11) = 4

To solve the equation, you must add 11 to each side of the equation:
x+(-11) + 11 = 4 + 11
x = 15

R^2 at SCC